# Self-supervised Learning of Event-guided Video Frame Interpolation for Rolling Shutter Frames

### v1

@@ 4.2 Latent Global Shutter Frame Generation
This module aims to generate a series of latent GS frames based on RS frames and DF, as shown in Fig 2 (II).Based on the analysis in Sec. 4.1, it can be inferred that the value of any point in the video can be determined by the value of a known point and the corresponding transformation in the $\mathcal{D}$.Therefore, masks are proposed to select the corresponding displacement field information for the reciprocal reconstruction between RS and GS frames, as shown in Fig. 4. The shape of the mask is $T\times H\times W$, and each item of the mask is a weight of motion for the corresponding position in DF. Specifically, we represent RS or GS frames as planes determined by their exposure time. We sample uniformly inside each item in the mask, and for each sampled point we calculate its direction to the RS and GS planes, as shown in Fig.4.For example, Fig. 4 (a) shows the mask $\mathcal{M}_{{r_{0}}\to{g_{0}}}$ of RS frame $I_{r_{0}}$ to GS frame $I_{g_{0}}$. Blue and red indicate the forward and backward deformation from $I_{r_{0}}$ to generate $I_{g_{0}}$.
We denote the mask that converts the $i$-th RS frame $I_{r_{i}}$ into the $j$-th GS frame $I_{g_{j}}$ as $\mathcal{M}_{{r_{i}}\to{g_{j}}}$. For the generation of GS frames, we use the Eq. 4 and Eq. 5, where $f_{warp}$ is the warping function  [13], $F_{{r_{0}}\to{g_{j}}}$ and $F_{{r_{1}}\to{g_{j}}}$ are warp weights for $I_{r_{0}}$ and $I_{r_{1}}$ to $I_{g_{j}}$, $I_{{r_{i}}\to{g_{j}}}$ is the $j$-th GS frame reconstructed from $i$-th RS frame, and $\otimes$ denotes the dot product and sum in the temporal dimension.
$\begin{split}F_{{r_{0}}\to{g_{j}}}=\mathcal{D}\otimes\mathcal{M}_{{r_{0}}\to{g_{j}}},F_{{r_{1}}\to{g_{j}}}=\mathcal{D}\otimes\mathcal{M}_{{r_{1}}\to{g_{j}}},\end{split}$ (4)
$\begin{split}I_{{r_{i}}\to{g_{j}}}=f_{warp}(I_{r_{i}},F_{r_{i}\to g_{j}}),i=\{0,1\}\end{split}$ (5)
@@
where $\Delta F_{{r_{0}}\to{g_{j}}},\Delta F_{{r_{1}}\to{g_{j}}}$ are residuals of the warp weights $F_{{r_{0}}\to{g_{j}}}$ and $F_{{r_{1}}\to{g_{j}}}$, and $O_{g_{j}}\in[0,1]$ is the degree of occlusion for $I_{r_{0}}$. Therefore, the refined latent $j$-th GS frame is generated, as formulated in Eq. 7:
@@
4.3 Reciprocal Reconstruction
As the ground truth GS frames are not available and the mapping from the input RS frames to latent GS frames is highly under-constrained, this module exploits how to reconstruct RS frames from a single GS frame for the purpose of self-supervision.As the displacement field includes the non-linear dense 3D spatiotemporal information of all pixels during the exposure time, we can simply achieve this target by Eq. 8 and Eq. 9, where $\mathcal{M}_{{g_{j}}\to{r_{i}}}$ is the mask from $j$-th GS frame $I_{g_{j}}$ to $i$-th RS frame $I_{r_{i}}$. Based on the definition of mask in Sec. 4.1, it indicates that masks of RS-to-GS and GS-to-RS are reversible for each other, namely $\mathcal{M}_{{g_{j}}\to{r_{i}}}=-\mathcal{M}_{{r_{i}}\to{g_{j}}}$.Because we can reconstruct RS frames $I_{r_{0}}$ or $I_{r_{1}}$ from each predicted latent GS frame $I_{g_{j}}$, we convert the supervision from the reconstruction of latent GS frames to the reconstruction of input RS frame.
$\begin{split}F_{{g_{j}}\to{r_{0}}}&=\mathcal{D}\otimes\mathcal{M}_{{g_{j}}\to{r_{0}}},\\ F_{{g_{j}}\to{r_{1}}}&=\mathcal{D}\otimes\mathcal{M}_{{g_{j}}\to{r_{1}}}\end{split}$ (8)
@@ 4.3 Reciprocal Reconstruction
Figure 6: Visual comparison on Gev-RS dataset.
@@
Table 2: Quantitative results (PSNR / SSIM / LPIPS) of the proposed framework and other methods on the Gev-RS simulated dataset. For supervised methods, TL refers to TimeLens[38], EU refers to EvUnRoll[46]. For the unsupervised method, TR refers to TimeReplayer[15]. LPIPS is calculated by the lpipis library[44] with AlexNet pretrained model[21]. ’Event’ indicates whether events are used, and ’SSL’ indicates whether it is self-supervised.
(A) GS-to-RS Loss: To realize the self-supervision, we reconstruct the input RS frames $I_{r_{0}}$, $I_{r_{1}}$ from the generated $i$-th frame $I_{g_{i}}$. For simplicity, we use the Charbonnier loss $\mathcal{L}_{c}$  [22] as the GS-to-RS self-supervision loss, formulated as:
$\mathcal{L}_{gs2rs}=\frac{1}{2N}\sum_{i=1}^{N}\left(\mathcal{L}_{c}(I_{{g_{i}}\to{r_{0}}},I_{r_{0}})+\mathcal{L}_{c}(I_{{g_{i}}\to{r_{1}}},I_{r_{1}})\right),$ (10)
where $N$ is the number of predicted GS frames.
(B) RS-to-RS Loss: Since the two input RS frames have spatiotemporal coherence, we leverage it as the constraint for imposing additional self-supervision. We reconstruct $i$-th RS frame by warping $j$-th RS frame and displacement field as Eq. 11 and Eq. 12. Then, we employ the $\mathcal{L}_{c}$ as our RS-to-RS loss as Eq. 13.
$I_{{r_{1}}\to{r_{0}}}=f_{warp}\left(I_{r_{1}},(\mathcal{D}\otimes\mathcal{M}_{{r_{1}}\to{r_{0}}})\right),$ (11)
$I_{{r_{0}}\to{r_{1}}}=f_{warp}\left(I_{r_{0}},(\mathcal{D}\otimes\mathcal{M}_{{r_{1}}\to{rs_{0}}})\right),$ (12)
$\mathcal{L}_{rs2rs}=\mathcal{L}_{c}(I_{{r_{0}}\to{r_{1}}},I_{r_{1}})+\mathcal{L}_{c}(I_{{r_{1}}\to{r_{0}}},I_{r_{0}})$ (13)
$\mathcal{L}=\lambda_{f}\mathcal{L}_{field}+\lambda_{rs}\mathcal{L}_{rs2rs}+\lambda_{gs}\mathcal{L}_{gs2rs}$ (14)
5 Experiments
@@
Implementation Details: We employ the Adam optimizer [20] for all experiments, with learning rates of $41e-$ 4 for Gev-RS, Fastec-RS and E-RS-VFI datasets, respectively. Our framework is trained for 100 epochs with the batch size of 4 using an NVIDIA RTX A30 GPU.
Evaluation Metrics: We evaluate our approach using the peak-signal-to-noise ratio (PSNR) [17] and structural similarity (SSIM) [40] and perceptual similarity metric LPIPS [44].
Figure 7: Visual comparison on Fastec-RS dataset.
@@
Dataset: 1) Gev-RS dataset [46] provides original videos captured by GS high-speed cameras with $1280\times 720$ resolution at 5700 fps. Notably, EvUnroll [46] considers both RS correction and deblurring, thus their simulation dataset includes blurry frames and lacks high frame rate frames for VFI evaluation.Therefore, we reconstruct RS frames and events from original videos to avoid the influence of blurring. Because tackling deblurring, RS correction, and VFI is extremely challenging. We first downsample the original videos to the same resolution($260\times 346$) of the DAVIS346 event camera [35]. Then, we input the resized frames to the event simulator vid2e [10] to synthesize event streams.We generate RS frames based on the same RS simulation process of Fastec-RS [24]. We follow the same dataset split scheme as in EvUnroll [46]: 20 videos for training and 9 videos for testing.2) Fastec-RS dataset  Fastec-RS dataset [24] provides the original frame sequences which are recorded by the high-speed GS cameras with the resolution of $640\times 480$ at 2400 fps. We follow the aforementioned settings to downsample frame sequences and generate the corresponding event streams and RS frames. Besides, we employ the same dataset split scheme as in Fastec-RS [24]: 56 sequences for training and 20 sequences for testing. 3) ERS-VFI dataset To evaluate our method on the real-world dataset, we use an ALPIX-Eiger event camera  [^1] https://www.alpsentek.com/product to collect a new dataset, called ERS-VFI. This camera outputs RGB frames with the resolution of $3264\times 2448$ and events with the resolution of $1632\times 1224$.For all collected videos, 19 videos are selected for training, and 10 videos are for testing. Finally, 3630 and 2071 frames with aligned events are used as the training and testing sets, respectively. To prevent memory overflow during training, we apply data augmentation strategies of [5], such as random crop, to our ERS-VFI dataset for all compared methods.
Table 3: Quantitative results (PSNR / SSIM / LPIPS) of the proposed framework and other methods on the Fastec-RS[24] dataset.
5.1 Comparison with SoTA Methods
We compare our method with five SoTA methods under three VFI settings: (I) one SoTA event-guided VFI method—TimeLens (TL) [38], which is based on GS frames. (II) Combined methods, event-guided RS correction method—EvUnRoll(EU) [46] + event-guided VFI methods: TL[38] or TimeReplayer (TR) [15]. (III) Frame-based RS-to-GS VFI methods: CVR [8] and RSSR [5]. Except for the frame-based settings (III) which only takes RS frames, the inputs of (I) and (II) are both RS frames and events, and the ground truth frames of these three settings are the GS frames. Evaluation on Gev-RS dataset:Tab. 2 presents the quantitative results from 4×4$\times$4 × to 32×32$\times$32 × interpolation, and the comparison of visual quality is shown in Fig. 6. Our method clearly outperforms the CVR and RSSR, which are frame-based RS-to-GS supervised VFI methods, in 4×4$\times$4 × interpolation by up to 1.32 dB in PSNR. In addition, our method has the best LPIPS scores among all the compared methods (expect 8×8$\times$8 × interpolation). Because our results do not suffer from the black holes, as shown in the red box of Fig. 6(g). This could be attributed to the capacity of estimating the complex non-linear motion of displacement field by utilizing the high temporal resolution of events.Without considering RS correction, the performances of TimeLens and the combination of TimeLens + EvUnroll degrade greatly with RS frames as inputs, and the interpolated results can not be recovered by RS correction methods.
Although the combination of two supervised methods —EvUnroll + TimeLens shows promising results owing to the high temporal resolution of events, such a two-step approach suffers from prohibitive computation costs (See Tab. 2). The performance of the combination of supervised method —EvUnroll + unsupervised method—TimeReplayer is limited, which demonstrates that the unsupervised VFI method can not be naively combined with RS correction methods.
@@ 5.1 Comparison with SoTA Methods
Evaluation on Fastec-RS dataset Here, we evaluate our methods on the Fastec-RS dataset, and the quantitative result is summarized in Tab. 3. We draw a similar conclusion as the experiments on the Gev-RS dataset, based on the quantitative comparison. Fig. 7 shows the visualization results of a outside street view captured by a moving camera and we have the sharpest reconstruction, for example, the number on the road sign (in the red box).
Evaluation on ERS dataset Fig. Self-supervised Learning of Event-guided Video Frame Interpolation for Rolling Shutter Frames shows the results of our method on the real-world dataset with 32$\times$ frame interpolation, and our method can generate successive smooth intermediate GS frames with spatialtemporal coherence, as shown in Fig. Self-supervised Learning of Event-guided Video Frame Interpolation for Rolling Shutter Frames (e). Also, Fig. 8 shows the GS frame (a) we reconstructed and the RS frame (b) we collected at time 0.25. By comparing the edge in the RS frames with that of events, we successfully correct the distorted edges, as shown in Fig.8 (a).
5.2 Ablation Studies
We conduct the following ablation experiments to evaluate the effectiveness of our proposed modules in Gev-RS dataset [46]: E-RAFT Channel, Time bins and iters of E-RAFT are set to 15, 6, and 12 in the baseline case which uses pretrained E-RAFT model [12]. Loss function: Tab.5 for each loss in Eq. 14. The 1st row represents the baseline, where solely the $\mathcal{L}_{gs2rs}$ loss (Eq.10) is utilized, In the 2nd and 3rd rows, the displacement field loss (Eq.3) and RS-to-RS loss (Eq.13) are incorporated concurrently. As a result, increases are observed in the PSNR (+0.12) and SSIM (+0.0015). Also, when perceptual loss is added in our self-supervised setting (Tab. 4), we find it hard to obtain favorable outcomes.
Time bins: We conduct the experiments to evaluate the influences of different time bins on the displacement field estimation. As shown in Tab. 4 (row 1-4), the settings with six time bins obtains the best PSNR and SSIM score. The explanation is that when the number of time bins goes up, the 3D deformation field will be divided into more segments bringing better performance because of more capacity of the non-linear motion. However, an excessive number of time bins can result in accumulated errors of motion and the degradation of performance. Fig. 9 shows warp weights corresponding to GS at different times with six time bins. It can be observed that the nonlinear motion of the bus is estimated clearly.
Perceptual loss: Perceptual loss $\mathcal{L}_{p}$ is widely used in previous work[15, 38, 18, 5, 8]. We also perform experiments to study the effectiveness of the perceptual loss [19] by applying the perpetual loss on GS-to-RS reconstruction and RS-to-RS warping. Tab. 4 (row 6-7) validates that introducing the perceptual loss can not boost the performance. Especially, applying perceptual loss only on the GS-to-RS reconstruction leads to the collapse of the displacement field and a worse performance. We argue that although the perceptual loss is applied in the compared method, it can not serve as the constraint in our self-supervised method.
Pretrained optical flow model: We evaluate the effect of the pre-train optical flow model by removing E-RAFT pre-train model.As shown in Tab. 4 (row 1,5), the experiments with pre-train E-RAFT model and without it demonstrate a similar performance. This indicates that our method is able to learn the optical flow without supervision.
Sample points of Height and Time of Mask: We investigate the impact of different sample points in height and time dimension during the estimation of the mask, as shown in Fig. 4. Tab. 4 (row 1,8-9) shows that the settings with different sample points have similar performance. This is because, for the setting with five height samples points and ten time samples points, each element has up to 50 samples points; thus, each element has ample samples to estimate the weight of warping.
Figure 8: (a) The reconstructed GS frame with $t=0.25$ in real-world dataset and events around $t=0.25$. (b) The RS frame and events around $t=0.25$. The green lines mark the edge of events and the edge of the RS frame which has obvious distortion.
Figure 9: The rolling shutter frames (a) and (e) and warp weights visualization. $F_{r_{i}\to g_{t}}$ means the warp weight form $i$-th RS frame to a GS frame at time $t$.
$T$ P $\mathcal{L}_{p}$ GS-to-RS $\mathcal{L}_{p}$ RS-to-RS $S_{h}$ $S_{t}$ PSNR SSIM
1 6 50 100 23.91 0.8091
2 4 50 100 23.78 0.8019
3 2 50 100 23.32 0.7928
4 12 50 100 23.67 0.7960
5 6 50 100 23.88 0.8082
6 6 50 100 18.50 0.6225
7 6 50 100 23.84 0.8134
8 6 5 10 23.90 0.8097
9 6 100 200 24.00 0.8109

### v2

@@ IV-B Latent Global Shutter Frame Generation
This module aims to generate a series of latent GS frames based on RS frames and DF, as shown in Fig 2 (II). Based on the analysis in Sec. IV-A, it can be inferred that the value of any point in the video can be determined by the value of a known point and the corresponding transformation in the $\mathbf{D}$. Therefore, weight maps are proposed to select the corresponding displacement field information for the reciprocal reconstruction between RS and GS frames, as shown in Fig. 4. The shape of the weight map is $T\times H\times W$, and each item of the weight map is a weight of motion for the corresponding position in DF. Specifically, we represent RS or GS frames as planes determined by their exposure time. We sample uniformly inside each item in the weight map, and for each sampled point we calculate its direction to the RS and GS planes, as shown in Fig.4. For example, Fig. 4 (a) shows the weight map $\mathbf{M}_{{r_{0}}\to{g_{0}}}$ of RS frame $I_{r_{0}}$ to GS frame $I_{g_{0}}$. Blue and red indicate the forward and backward deformation from $I_{r_{0}}$ to generate $I_{g_{0}}$.
As illustrated in Fig. 5, which depicts the schematic diagram of the entire process, we define the weight map that converts the $i$-th RS frame $I_{r_{i}}$ into the $j$-th GS frame $I_{g_{j}}$ as $\mathbf{M}_{{r_{i}}\to{g_{j}}}$. For the generation of GS frames, we use Eq. 4 and Eq. 5, where $f_{\text{warp}}$ is the warping function [64]. $F_{{r_{0}}\to{g_{j}}}$ and $F_{{r_{1}}\to{g_{j}}}$ are the estimated flow fields used to warp the RS frames $I_{r_{0}}$ and $I_{r_{1}}$ toward the target GS frame $I_{g_{j}}$, and $I_{{r_{i}}\to{g_{j}}}$ denotes the reconstructed GS frame from the $i$-th RS frame. The warp fields $F_{{r_{i}}\to{g_{j}}}$ are obtained by combining the predicted displacement field $\mathbf{D}$ and a learned temporal weight map $\mathbf{M}$ through a weighted summation along the temporal axis. The operation is defined as Eq. 4.
$\begin{split}F_{{r_{0}}\to{g_{j}}}=\mathbf{D}\otimes\mathbf{M}_{{r_{0}}\to{g_{j}}},\quad F_{{r_{1}}\to{g_{j}}}=\mathbf{D}\otimes\mathbf{M}_{{r_{1}}\to{g_{j}}},\end{split}$ (4)
Here, $\mathbf{D}\in R^{2\times T\times H\times W}$ is the displacement field, representing motion vectors in both horizontal and vertical directions over $T$ temporal bins. The weight maps $\mathbf{M}_{{r_{i}}\to{g_{j}}}\in R^{T\times H\times W}$ assign importance to each displacement sample. The operation $\otimes$ performs element-wise multiplication between the two tensors followed by summation along the temporal dimension $T$, yielding a final flow field $F\in R^{2\times H\times W}$. Finally, each warped GS frame is obtained using the warping function as Eq. 5. This process enables temporally and geometrically consistent synthesis of global shutter images from temporally misaligned rolling shutter inputs.
$\begin{split}I_{{r_{i}}\to{g_{j}}}=f_{\text{warp}}(I_{r_{i}},F_{{r_{i}}\to{g_{j}}}),\quad i\in\{0,1\},\end{split}$ (5)
@@
where $\Delta F_{{r_{0}}\to{g_{j}}},\Delta F_{{r_{1}}\to{g_{j}}}$ are residuals of the warp weights $F_{{r_{0}}\to{g_{j}}}$ and $F_{{r_{1}}\to{g_{j}}}$, and $O_{g_{j}}\in[0,1]$ is the degree of occlusion for $I_{r_{0}}$. Therefore, the refined latent $j$-th GS frame is generated, as Eq. 7:
@@
IV-C Reciprocal Reconstruction
As the ground truth GS frames are not available and the mapping from the input RS frames to latent GS frames is highly under-constrained, this module exploits how to reconstruct RS frames from a single GS frame for the purpose of self-supervision. As the displacement field includes the non-linear dense 3D spatiotemporal information of all pixels during the exposure time, we can simply achieve this target by Eq. 8 and Eq. 9, where $\mathbf{M}_{{g_{j}}\to{r_{i}}}$ is the weight map from $j$-th GS frame $I_{g_{j}}$ to $i$-th RS frame $I_{r_{i}}$. Based on the definition of weight map in Sec. IV-A, it indicates that weight maps of RS-to-GS and GS-to-RS are reversible for each other, namely $\mathbf{M}_{{g_{j}}\to{r_{i}}}=-\mathbf{M}_{{r_{i}}\to{g_{j}}}$. Because we can reconstruct RS frames $I_{r_{0}}$ or $I_{r_{1}}$ from each predicted latent GS frame $I_{g_{j}}$, we convert the supervision from the reconstruction of latent GS frames to the reconstruction of input RS frame.
$\begin{split}F_{{g_{j}}\to{r_{0}}}&=\mathbf{D}\otimes\mathbf{M}_{{g_{j}}\to{r_{0}}},\\ F_{{g_{j}}\to{r_{1}}}&=\mathbf{D}\otimes\mathbf{M}_{{g_{j}}\to{r_{1}}}\end{split}$ (8)
@@ IV-C Reciprocal Reconstruction
(A) GS-to-RS Loss: To realize the self-supervision, we reconstruct the input RS frames $I_{r_{0}}$, $I_{r_{1}}$ from the generated $i$-th frame $I_{g_{i}}$. For simplicity, we use the Charbonnier loss $\mathbf{L}_{c}$  [65] as the GS-to-RS self-supervision loss, formulated as:
$\mathbf{L}_{gs2rs}=\frac{1}{2N}\sum_{i=1}^{N}\left(\mathbf{L}_{c}(I_{{g_{i}}\to{r_{0}}},I_{r_{0}})+\mathbf{L}_{c}(I_{{g_{i}}\to{r_{1}}},I_{r_{1}})\right),$ (10)
where $N$ is the number of predicted GS frames.
Figure 7: Visual comparison on Fastec-RS dataset [36], where two random time points are selected as target frames for recovery. The results of RSSR (g) exhibit black holes, particularly evident in scene (II). While TimeLens (c) and TimeLens+EvUnRoll (d) show noticeable distortion and artifacts when rolling shutter distortion is not accounted for. Our results demonstrate competitiveness with supervised methods (e.g., CVR and EvUnRoll + TimeLens), and even outperform them in accurately reconstructing the speed limit sign within scene (II).
TABLE II: Quantitative results (PSNR / SSIM / LPIPS) of the proposed framework and other methods on the Fastec-RS[36] dataset. Bold indicates the best performance.
(B) RS-to-RS Loss: Since the two input RS frames have spatiotemporal coherence, we leverage it as the constraint for imposing additional self-supervision. We reconstruct $i$-th RS frame by warping $j$-th RS frame and displacement field as Eq. 11 and Eq. 12. Then, we employ the $\mathbf{L}_{c}$ as our RS-to-RS loss as Eq. 13.
$I_{{r_{1}}\to{r_{0}}}=f_{warp}\left(I_{r_{1}},(\mathbf{D}\otimes\mathbf{M}_{{r_{1}}\to{r_{0}}})\right),$ (11)
$I_{{r_{0}}\to{r_{1}}}=f_{warp}\left(I_{r_{0}},(\mathbf{D}\otimes\mathbf{M}_{{r_{1}}\to{rs_{0}}})\right),$ (12)
$\mathbf{L}_{rs2rs}=\mathbf{L}_{c}(I_{{r_{0}}\to{r_{1}}},I_{r_{1}})+\mathbf{L}_{c}(I_{{r_{1}}\to{r_{0}}},I_{r_{0}})$ (13)
$\mathbf{L}=\lambda_{f}\mathbf{L}_{field}+\lambda_{rs}\mathbf{L}_{rs2rs}+\lambda_{gs}\mathbf{L}_{gs2rs}$ (14)
Figure 8: Visual comparison on Gev-RS dataset [4]. Similar to the results on the Fastec-RS dataset, RSSR’s results (g) exhibit black holes, whereas TimeLens (c) and TimeLens+EvUnRoll (d) show noticeable distortion and artifacts. Importantly, our results accurately reconstruct the details of the poster (highlighted in the red box area), especially the face that other supervised methods find challenging to accomplish.
TABLE III: Quantitative results (PSNR / SSIM / LPIPS) of the proposed framework and other methods on the Gev-RS simulated dataset. For supervised methods, TL refers to TimeLens[3], EU refers to EvUnRoll[4]. For the unsupervised method, TR refers to TimeReplayer[46]. LPIPS is calculated by the lpipis library[66] with AlexNet pretrained model[67]. ’Event’ indicates whether events are used, and ’SSL’ indicates whether it is self-supervised. Black bold indicates the best performance.
@@
V Experiments
@@
Implementation Details: We employ the Adam optimizer [68] for all experiments, with learning rates of $41e-$ 4 for all datasets. Our framework is trained for 100 epochs with a batch size of 4 using an NVIDIA RTX A30 GPU.
@@
Evaluation Metrics: We evaluate our approach using the peak-signal-to-noise ratio (PSNR) [69] and structural similarity (SSIM) [70] and perceptual similarity metric LPIPS [66]. Dataset: 1) Fastec-RS dataset  Fastec-RS dataset [36] provides the original frame sequences which are recorded by the high-speed GS cameras with the resolution of $640\times 480$ at 2400 fps. This dataset offers an external perspective of a fast-moving vehicle and can partially capture scenes resembling those of high-speed flying aircraft. We first downsample the original videos to the same resolution($260\times 346$) of the DAVIS346 event camera [41]. Then, we input the resized frames to the event simulator vid2e [71] to synthesize event streams. We generate RS frames based on the same RS simulation process of Fastec-RS [36]. Besides, we employ the same dataset split scheme as in Fastec-RS [36]: 56 sequences for training and 20 sequences for testing. 2) Gev-RS dataset [4] provides original videos captured by GS high-speed cameras with $1280\times 720$ resolution at 5700 fps. We follow the aforementioned settings to downsample frame sequences and generate the corresponding event streams and RS frames. We follow the same dataset split scheme as in EvUnroll [4]: 20 videos for training and nine videos for testing. Notably, EvUnroll [4] considers both RS correction and deblurring. Thus, their simulation dataset includes blurry frames and lacks high frame rate frames for VFI evaluation. Therefore, we reconstruct RS frames and events from original videos to avoid the influence of blurring. 3) ERS dataset To evaluate our method on the real-world dataset, we use an ALPIX-Eiger event camera [1] to collect a new dataset called ERS. This camera outputs RGB frames with the resolution of $3264\times 2448$ and events with the resolution of $1632\times 1224$. For all collected videos, 19 videos are selected for training, and ten videos are for testing. Finally, 3630 and 2071 frames with aligned events are used as the training and testing sets, respectively. To prevent memory overflow during training, we apply data augmentation strategies of [39], such as random crop, to our ERS dataset for all compared methods. 4) UAV-RS dataset The first two datasets mentioned above have been widely adopted as benchmarks for quantitative evaluations in the academic community. In contrast, our third dataset is designed to facilitate real-world qualitative assessments. To comprehensively evaluate the robustness and generalizability of our method in various aerial photography scenarios, we introduced a new dataset based on aerial photography. We deployed a DJI drone to capture nine high-frame-rate videos with $120fps$ frame rate, from which we generated simulated events and RS frames. The resolution of the generated rolling shutter frame is 260 x 346, with a frame rate of 0.46 fps (calculated as 120/260). This allowed us to conduct a qualitative evaluation of our method’s effectiveness in aerial imaging contexts.
V-A Comparison with SoTA Methods
We compare our method with five SoTA methods under three VFI settings: one SoTA event-guided VFI method—TimeLens (TL) [3], which is based on GS frames. Combined methods, event-guided RS correction method—EvUnRoll (EU) [4] + event-guided VFI methods: TL [3] or TimeReplayer (TR) [46]. Frame-based RS-to-GS VFI methods: CVR [17] and RSSR [39]. Except for the frame-based methods which only takes RS frames, the inputs of other methods are both RS frames and events, and the ground truth frames of these three settings are the GS frames.
@@ V-A Comparison with SoTA Methods
Evaluation on Fastec-RS dataset: We evaluate our methods on the Fastec-RS dataset, and the quantitative result is summarized in Tab. II. We draw a similar conclusion as the experiments on the Gev-RS dataset, based on the quantitative comparison. Fig. 7 shows the visualization results of an outside street view captured by a moving camera and we have the sharpest reconstruction, for example, the number on the road sign (in the red box). For more visualization results, please refer to the supplementary material (Faster.zip).
Evaluation on ERS dataset By comparing the edge in the RS frames with that of events, we successfully correct the distorted edges, as shown in Fig.9 (a). For more visualization results, please refer to the supplementary material (DJI.zip).
Evaluation on Gev-RS dataset: Tab. III presents the quantitative results from 4×4$\times$4 × to 32×32$\times$32 × interpolation, and the comparison of visual quality is shown in Fig. 8. Our method clearly outperforms the CVR and RSSR, which are frame-based RS-to-GS supervised VFI methods, in 4×4$\times$4 × interpolation by up to 1.32 dB in PSNR. In addition, our method has the best LPIPS scores among all the compared methods (expect 8×8$\times$8 × interpolation). Because our results do not suffer from the black holes, as shown in the red box of Fig. 8 (g). This could be attributed to the capacity to estimate the complex non-linear motion of the displacement field by utilizing the high temporal resolution of events. For more visualization results, please refer to the supplementary material (Gev-RSC.mp4).
V-B Efficiency Evaluation
In this section, we delve into an evaluation of our method, focusing primarily on its generalization, inference speed, transmission bandwidth, etc. This analysis aims to provide a comprehensive understanding of the efficiency of our model and its robustness.
Generalization testing on UAV-RS dataset: In Fig. 6, we present the visualization of our model on the UAV-RS dataset. Having been trained on the Fastec dataset, the model’s evaluation on the UAV dataset serves to underscore its generalization capabilities. Impressively, our approach adeptly rectifies the RS deformation, and the outcomes of the frame interpolation are notably effective.
Figure 9: Visualization results from real data: (I) showcases the input, consisting of events and two successive RS frames. Following RS correction, the output is presented in (II). By overlaying the events onto both our output and the initial input, it becomes evident that our method adeptly rectifies RS deformations. The interpolated frame output is depicted in (III), where our technique achieves a 32-fold interpolation.The green dashed boxes highlight temporal artifacts that appear in competing methods. Please zoom in to observe these details more clearly.